# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ 4. Proof of Main Theorem
(4.32) ${\big{\|}\tilde{h}\big{\|}}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}},$
${\big{\|}\tilde{h}\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
On the other hand, we know
@@
Combining (4.25), (4.29), (4.31), (4.32) with (4.33), we obtain
(4.34) ${\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\qquad\cdot\big{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\qquad\cdot\big{(}{\big{\|}(\omega_{1},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}(\omega_{2},A_{2})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}.$
This implies the uniqueness of solution $(\omega,A)$ in ${\mathcal{X}}_{\alpha,M}$ to (1.15) when ${\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{\delta_{0}}{M^{\frac{5+3\varepsilon}{2}}}$ with $\delta_{0}=\delta_{0}(\alpha,\varepsilon)\ll 1$.
Appendix
This appendix provides a brief overview of the triple-deck formalism, based on the physical model of an incompressible boundary layer with a roughness element; for further details, see [26]. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and its streamwise length scale is $O(\epsilon^{3})$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (near $X=1$). The three distinct layers are summarized as follows:
I. Viscous sublayer: Vertical scale $O(\epsilon^{5})$, streamwise scale $O(\epsilon^{3})$.
II. Main layer: Vertical scale $O(\epsilon^{4})$, streamwise scale $O(\epsilon^{3})$.
III. Upper layer: Vertical scale $O(\epsilon^{3})$, streamwise scale $O(\epsilon^{3})$.
@@
$x_{*}=\epsilon^{-3}(X-1)=O(1),\quad\tilde{y}=\epsilon^{-4}Y=O(1).$
The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+O(\epsilon^{4})$ and a mean-flow distortion,
@@
The solutions of this system are
@@
$y_{*}=\epsilon^{-5}Y=O(1).$
Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$. Thus, the main-layer solution behaves like $u_{NS}=\bar{U}_{0}+\epsilon\tilde{u}+o(\epsilon)\to\epsilon(\lambda y_{*}+A_{*})+o(\epsilon)$, which determines the magnitude of the sublayer-layer solution. From the balance of the NS equations, we can expand the solution as,
@@
The upper boundary condition of the sublayer equations should be obtained by matching with the main layer. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Then for any $\beta_{0}>0$, we have
@@
From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as following:
@@
Indeed, the above nonlinear system is usually solved using numerical approach. In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$. Consequently, the solution $\phi_{*}=(u_{*},v_{*},P_{*})$ should be denoted by $\phi_{*,\epsilon}$ to include the impact of the upper boundary. It is clear that $\phi_{*,\epsilon}$ also satisfy the system (A.17), except for the upper boundary condition (A.16), which is converted to
@@
Acknowledgments. M. Dong is partially supported by the Chinese Academy of Sciences (CAS) Project for Young Scientists in Basic Research under Grant No. YSBR-087. C. Wang is partially supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12471189. Z. Zhang is partially supported by the NSFC under Grant No. 12288101.

### v2

@@ 4. Existence and uniqueness of classical solution
(4.17) ${\big\|\tilde{h}\big\|}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}},$
${\big\|\tilde{h}\big\|}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}.$
On the other hand, we know that
@@
Summing up (4.13), (4.15), (4.16), (4.17) and (4.18), we obtain
${\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}$
$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big)$
$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|(\omega_{1},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|(\omega_{2},A_{2})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big).$
This implies the uniqueness of solution $(\omega,A)$ in ${\mathcal{X}}_{\alpha,M}$ to (1.13) when ${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{\delta_{0}}{M^{\frac{5+3\varepsilon}{2}}}$ with $\delta_{0}=\delta_{0}(\alpha,\varepsilon)\ll 1$. ∎
Appendix A The physical derivation of triple-deck equation
This appendix provides a brief overview of the triple-deck formalism, based on the physical model of an incompressible boundary layer with a roughness element; for further details, see [26]. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be ${\mathcal{O}}(\epsilon^{5})$ or smaller, and its streamwise length scale is ${\mathcal{O}}(\epsilon^{3})$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (near $X=1$). The three distinct layers are summarized as follows.
I. Viscous sublayer: Vertical scale ${\mathcal{O}}(\epsilon^{5})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
II. Main layer: Vertical scale ${\mathcal{O}}(\epsilon^{4})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
III. Upper layer: Vertical scale ${\mathcal{O}}(\epsilon^{3})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
@@ Appendix A The physical derivation of triple-deck equation
$x_{*}=\epsilon^{-3}(X-1)={\mathcal{O}}(1),\quad\tilde{y}=\epsilon^{-4}Y={\mathcal{O}}(1).$
The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+{\mathcal{O}}(\epsilon^{4})$ and a mean-flow distortion,
@@
The solutions to this system are
@@
$y_{*}=\epsilon^{-5}Y={\mathcal{O}}(1).$
Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+{\mathcal{O}}(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$. Thus, the main-layer solution behaves like $u_{NS}=\bar{U}_{0}+\epsilon\tilde{u}+o(\epsilon)\to\epsilon(\lambda y_{*}+A_{*})+o(\epsilon)$, which determines the magnitude of the sublayer-layer solution. From the balance of the NS equations, we can expand the solution as,
@@
The upper boundary condition of the sublayer equations should be obtained by matching with the main layer. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}={\mathcal{O}}(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Then for any $\beta_{0}>0$, we have
@@
From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as follows
@@
Indeed, the above nonlinear system is usually solved using numerical approach. In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in{\mathcal{O}}(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$. Consequently, the solution $\phi_{*}=(u_{*},v_{*},P_{*})$ should be denoted by $\phi_{*,\epsilon}$ to include the impact of the upper boundary. It is clear that $\phi_{*,\epsilon}$ also satisfy the system (A.17), except for the upper boundary condition (A.16), which is converted to
@@
Acknowledgments
M. Dong is partially supported by the NSFC under Grant No. 12588201 and the CAS Project for Young Scientists in Basic Research under Grant No. YSBR-087. C. Wang is partially supported by the NSFC under Grant No. 12471189. Z. Zhang is partially supported by the NSFC under Grant No. 12288101.
Data availability The manuscript has no associated data.
Declarations
Conflict of interest The authors state that there is no conflict of interest.